1 a.) Find the area in between Z = -2.05 and Z =1.68.

Round to 4 decimal places

1 b.)

In a random sample of 519 judges, it was found that 285 were introverts. Calculate a 90% confidence interval for the proportion of all judges who were introverts.

The lower bound is and the upper bound is .

Round your bounds to 3 decimal places.
Answer #1

1a) P.(-2.05<z< 1.68) = P(z<1.68) -P(z<-2.05) = 0.9535 -0.0202 (using z value table) =0.9333Problem 1b) Given 285 519 0.5491 Proportion 0.9 0.1 Critical Value Z0/2 = 1.64485363 + 1.645 90 % Confidence interval is =ộ +